A right triangle has a hypotenuse of length 20 and a leg of length 12. What is the length of the other leg? If necessary, round your answer to two decimal places.
Answer by cleomenius(959)

a^2 + b^2 = c^2
12^2 + b^2 = 20^2
144 + b^2 = 400
b^2 = 256
16 is the length of the other leg.
Cleomenius.